### v1

# Breaking the Visual Shortcuts in Multimodal Knowledge-Based Visual Question Answering

Abstract
Existing Multimodal Knowledge-Based Visual Question Answering (MKB-VQA) benchmarks suffer from “visual shortcuts”, as the query image typically matches the primary subject entity of the target document. We demonstrate that models can exploit these shortcuts, achieving comparable results using visual cues alone. To address this, we introduce Relational Entity Text-Image kNowledge Augmented (RETINA) benchmark, automatically constructed using an LLM-driven pipeline, consisting of 120k training and 2k human-curated test set. RETINA contains queries referencing secondary subjects (i.e. related entities) and pairs them with images of these related entities, removing the visual shortcut. When evaluated on RETINA existing models show significantly degraded performance, confirming their reliance on the shortcut. Furthermore, we propose Multi-Image MultImodal Retriever (MIMIR), which enriches document embeddings by augmenting images of multiple related entities, effectively handling RETINA, unlike prior work that uses only a single image per document. Our experiments validate the limitations of existing benchmarks and demonstrate the effectiveness of RETINA and MIMIR. Our project is available at: Project Page.
1 Introduction
Knowledge-Based Visual Question Answering (KB-VQA) is a challenging task where not only the visual input and its corresponding natural language query are processed, but external world knowledge must be integrated and reasoned over to correctly determine the answer. While early KB-VQA approaches were reliant on unimodal, text-only KBs [chen2023can, mensink2023encyclopedic, marino2019okvqavisualquestionanswering, schwenk2022aokvqabenchmarkvisualquestion, lin2024preflmr, lin2023finegrainedlateinteractionmultimodalretrieval], recent research [deng-etal-2025-muka, caffagni2025recurrence, wei2024uniir, lin2025mmembeduniversalmultimodalretrieval, yu2025cafeunifyingrepresentationgeneration, liu2024lamralargemultimodalmodel, kim2025genius] is shifting to the development of Multimodal KB-VQA (MKB-VQA) to integrate richer information.
(a) QA in Previous MKB-VQA
(b) QA in RETINA
(c) GT Document
Figure 1: Comparison of Previous MKB-VQA Benchmarks and RETINA. (a) In previous benchmarks, the query image correspond to the same visual entity (highlighted in red) as (c) the shared GT document, creating a visual shortcut. (b) In RETINA, the query uses a different visual entity (highlighted in green), breaking the visual shortcut.
In the real world, information relevant to a query is frequently found not in documents directly about the query image but in other related documents. For instance, when asking “Which beetle, native to North America, feeds on this particular plant?” with a photo of an potato, the answer “Lema daturaphila” may be found not in the “Potato” document but rather in the “Lema daturaphila” document, as shown in Fig. 1b. As in such cases, it is often uncertain which document contains the correct answer. However, existing MKB-VQA benchmarks are typically limited to scenarios where each query’s image entity directly matches the primary subject (i.e., main entity) of the target document and its image, leading to a visual shortcut. As illustrated in  Fig. 1a, given a query “In which part of the world does this animal live?” accompanied by an image of the beetle “Lema daturaphila”, existing benchmarks allow trivial retrieval by exploiting visual cues that directly match the main entity. Consequently, models trained on such data tend to rely on these superficial visual correlations. To verify this shortcut, a preliminary experiment where a model [deng-etal-2025-muka] was trained without using the textual component of the query was conducted. As shown in Fig. 2, the model achieved comparable retrieval performance using only the query image, indicating that existing MKB-VQA benchmarks are largely solvable through visual shortcut exploitation.
To address this, we construct Relational Entity Text-Image kNowledge Augmented (RETINA) bench, a MKB-VQA benchmark designed to eliminate such shortcuts and better reflect real-world. RETINA is built through an automated, LLM-driven pipeline that (1) identifies entities related to each document’s main entity (i.e. related entity), (2) constructs a structured knowledge graph, (3) extracts subgraphs for query generation, and (4) generates queries guided by those subgraphs. As prompt-based approaches that feed full documents to LLMs often produce hallucinated or incorrect entity references, we instead employ a graph-guided generation process [ainslie2023gqa, tran2025reasonvqa, trivedi2022musiquemultihopquestionssinglehop, reddy-etal-2017-generating]. This strategy allows controlled query construction by extracting relevant subgraphs corresponding to specific related entities. The resulting queries are then paired with images of the corresponding related, but non-main entities, removing the visual shortcut. Our automated pipeline produces 120K generated training samples alongside a 2K human-verified test set. When evaluating existing models  [lin2024preflmr, deng-etal-2025-muka, wei2024uniir, caffagni2025recurrence], which were trained on shortcut-prone benchmarks, we observe a substantial performance drop on RETINA, reflecting the models’ design that typically augments each document with a single image of its main entity.
To overcome this limitation, we propose Multi-Image MultImodal Retriever (MIMIR) that enriches document representations by integrating images of multiple related entities, extending the single-image framework of prior work [deng-etal-2025-muka]. Specifically, MIMIR extracts related entities from document text, augments multiple images of the related entities, which guides the retrieval of the corresponding documents that contain the related entities, thereby enabling retrieval in the absence of visual shortcuts. We summarize our contributions as follows:
  • We identify limitations in existing MKB-VQA benchmarks and develop RETINA, a new benchmark that captures realistic scenarios without visual shortcuts.
  • We propose MIMIR, a retriever designed to address these realistic queries with or without visual shortcuts.
  • We demonstrate the effectiveness of the RETINA and MIMIR through experiments.
(a) Infoseek
(b) EVQA
Figure 2: Preliminary Experiment on Existing Benchmarks. Recall@$k$ on benchmarks such as Infoseek [chen2023can] and EVQA [mensink2023encyclopedic], which exhibit visual shortcuts [deng-etal-2025-muka]. Models fine-tuned with image-only (blue) queries versus image+text (green) queries.
2 Related Works
Knowledge-based VQA Benchmarks Early developments in KB-VQA [lerner2022viquae, lin2024preflmr, hu2023avis, lin2023fine, marino2019okvqavisualquestionanswering, schwenk2022aokvqabenchmarkvisualquestion, jain2021select, wang2017fvqa, shahMYP19] incorporated external textual KBs such as Wikipedia, enabling responses grounded in external knowledge. Later developments [deng-etal-2025-muka, caffagni2025recurrence, wei2024uniir] expanded the textual KBs into the multimodal KBs by associating each textual document with corresponding images. However, these existing MKB-VQA benchmarks are biased towards samples with visual shortcuts, where the query’s image depicts the same visual entity as the document’s image entity. To address this bias, we introduce RETINA, a new benchmark that explicitly eliminates such shortcuts.
Multimodal Retrievers Early multimodal retrievers  [radford2021learning, Mori1999ImagetowordTB, kolesnikov2020bigtransferbitgeneral] primarily focused on image-to-text and text-to-image retrieval by aligning visual and textual representations within a shared embedding space. Subsequent studies  [Gao_2022_CVPR, lin2023finegrainedlateinteractionmultimodalretrieval, lin2024preflmr, long2025retrieval, lin2022revive] extended this paradigm to handle queries that jointly incorporate both textual and visual inputs. Recent multimodal KB retrieval approaches [deng-etal-2025-muka, caffagni2025recurrence] represent documents with both textual and visual modalities and demonstrate strong performance on tasks involving fine-grained, knowledge-intensive entities. In parallel, universal retrieval frameworks have been proposed [wei2024uniir, yu2025cafeunifyingrepresentationgeneration, lin2025mmembeduniversalmultimodalretrieval, kim2025genius, liu2025lamra] enabling retrieval across arbitrary query and document modality pairs. These existing retrievers encode each document with a single image, but real-world queries often involve diverse entities. Our MIMIR employs multi-image encoding to better handle realistic multimodal queries.
3 Dataset Construction
Figure 3: RETINA Benchmark Generation Pipeline. (a) constructs one-hop neighborhood graphs by extracting named entities (gray box) related to the answer entity (white box) and their relations using an LLM [bai2023qwen]; (b) samples an query entity (green) and qualifying entity (teal) to form a target subgraph with the answer entity (red); and (c) feeds the target subgraph into an LLM to generate a textual query and collect a corresponding image from M2KR Images [lin2024preflmr]. The query is then paraphrased to minimize lexical overlap with the document.
Our goal is to build an MKB-VQA benchmark that captures challenging real-world scenarios, specifically those where visual shortcuts are absent. In practice, information associated with a particular entity is often scattered across multiple documents rather than confined to the one describing that entity itself. Consequently, user are inherently uncertain about which document contains the answer to a given query. However, existing MKB-VQA benchmarks are biased toward scenarios where the main entity of the target document typically has the same entity as the query image. To address this bias, we focus on collecting samples that do not permit such visual shortcuts, by deliberately selecting a query image that differ from the main entity of the target Wikipedia document. Specifically, we employ an LLM [bai2023qwen] to identify entities related to the main entity within the target document, and generate a multimodal query by selecting the query image from these related entities. Through this automatic generation pipeline, we construct a large-scale benchmark RETINA, consisting of 120K training samples and a 2K human-curated test set. When combined with existing benchmarks, our RETINA bench provides a realistic evaluation setting for MKB-VQA in real-world scenarios.
3.1 MKB-VQA Sample Generation Pipeline
Automatically generating VQA samples from raw Wikipedia documents is challenging for LLMs because the documents are often long and the relationships between entities are implicit and scattered throughout the text [liu2023lostmiddlelanguagemodels]. We address this by explicitly extracting relationships between entities from the documents before generating VQA samples. Specifically, we adopt a three-stage pipeline: (a) converting the document into a knowledge graph to enable better control over the generation process, (b) selecting target subgraphs, and (c) generating multimodal QA based on the selected subgraphs.
Knowledge Graph Construction Given the text body of a Wikipedia document and its corresponding main entity, we first construct a one-hop neighborhood graph in which the connected nodes represent named entities that appear in direct relation to the main entity. To obtain these entities, we prompt an LLM to extract all named entities that explicitly co-occur with the main entity and exhibit a semantic relationship to it. Note that we do not fully convert the document into a structured entity–relation graph, as doing so often discards important contextual information present in natural language, such as causal descriptions, attributes, temporal dependencies, and explanatory qualifiers. Instead, we build a one-hop graph in which each related entity is tagged with a natural language sentence that captures its relation to the main entity while preserving the original descriptive details. This representation allows precise control over which entities to select as targets for QA generation, while preserving access to the rich contextual nuance and descriptive detail encoded in the original sentences. Figure 3a) illustrates this process. For the Wikipedia page on ‘Lema daturaphila’, the LLM identifies related entities such as ‘Potato’ and ‘Solanaceae’, producing a one-hop graph in which these entities appear as neighboring nodes.
Target Subgraph Extraction MKB-VQA samples are designed to involve two distinct visual–semantic concepts: one depicted by the query image and another serving as the answer. As we focus on question answering grounded in the Wikipedia KB, we represent these concepts using two entities drawn from each document. To avoid visual shortcuts, the document’s main entity cannot serve as the concept depicted in the query image. We therefore assign the main entity as the answer entity and randomly select one of its related entities from the constructed graph as the query entity. Each such pair, together with its tagged sentence describing their relation, forms the foundation for generating an MKB-VQA sample. However, directly formulating a question from this pair can lead to answer ambiguity, where multiple entities could validly satisfy the question. For example, as shown in Fig. 3b, the graph includes a relation between ‘Lema daturaphila’ and ‘Potato’ expressed as ‘Lema daturaphila feeds on potato plants, though less frequently than on tomatillos.’ Using an image of Potato as the query may produce the question ‘What beetle feeds on this plant?’, which fails to uniquely identify Lema daturaphila as the answer. To mitigate this answer ambiguity, we introduce a qualifying entity—an additional related entity that provides contextual cues to uniquely specify the answer entity among multiple possible candidates. This design substantially reduces answer ambiguity, ensuring that most generated questions yield a single, well-defined answer while preserving the richness and diversity of real-world multimodal reasoning scenarios. As a result, we extract the two-hop subgraph containing the query, answer, and qualifying entities, and pass it to the LLM to generate the QA surface form.
LLM-based VQA Sample Generation Given an extracted target subgraph, our goal is to generate a QA pair for MKB-VQA. We first obtain a query image by retrieving an image of the query entity from M2KR [lin2024preflmr]. Then, we prompt the LLM to formulate a question whose answer should correspond to the answer entity, conditioned on two natural language descriptions describing the relations among the entities in the subgraph. To prevent textual leakage, we explicitly instruct the LLM not to mention the query entity by name within the question, but instead to refer to it using an appropriate demonstrative phrase (e.g., “this organization” or “this monument”), ensuring that the model must infer it from the query image rather than the text. Finally, the generated question is passed to the LLM once more for paraphrasing, and we filter out questions that can be retrieved by BM25 [robertson1995okapi] in the top-5 results, thereby reducing lexical overlap with the source document. The full prompts are provided in the Supp. Mat., and Fig. 3c) illustrates this process along with an example QA pair.
3.2 RETINA Bench Construction
Following the procedures described above, we construct the RETINA Bench. It comprises a large-scale training set with 120k samples and a 2k test set. The test set is manually curated by human experts to ensure factual accuracy, linguistic fluency, and answer unambiguity (see Supp. Mat. for details). We divide the test set into seen and unseen subsets. The seen subset refers to samples whose GT documents overlap with the training data, while the unseen subset contains samples without such overlap. To ensure a single GT document per query, we filter out from each one-hop graph any related entity whose corresponding document (where it serves as the main entity) also lists the original main entity as a related entity. This prevents cases where both the main entity’s document and a related entity’s document would reveal the answer.
Figure 4: Overview of MIMIR Document Encoder Architecture. (a) given a document, related named entities are identified with an LLM [bai2023qwen], and corresponding images are collected from the KB; (b) textual, global image, and patch-level features are extracted, with patch features attending to textual features through cross-attention to yield multimodal features; and (c) entity token embeddings are incorporated into the textual features prior to cross-attention for richer contextualization; (d) the final document embedding jointly integrates textual, global, and multimodal features projected into the same embedding space.
4 Method
Existing multimodal retrievers [deng-etal-2025-muka, caffagni2025recurrence, wei2024uniir] typically encode each document together with an image depicting its main entity. While these approaches achieve strong performance on existing MKB-VQA benchmarks, their effectiveness diminishes on RETINA. This limitation arises because each document embedding is augmented solely with the image of its main entity. Consequently, these systems often fail to retrieve the correct document when the query image depicts a non-main entity within the target document. To address this issue, we introduce MIMIR, a multimodal retriever that enriches document representations by incorporating images of multiple related entities. Our approach builds upon MuKA [deng-etal-2025-muka], a multimodal single-image document retriever, and extends its representation capacity. In the following, we first review the scoring mechanism and query encoder of [deng-etal-2025-muka] (Sec. 4.1 and 4.2), and then describe our extended document embedding module in detail (Sec. 4.3).
4.1 Late Interaction Scoring
Following ColBERT family of multimodal retrievers [deng-etal-2025-muka, lin2024preflmr], we represent both the query and document as sets of token-level vector, $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{m}\}$ and $\mathcal{D}=\{d_{1},d_{2},\dots,d_{n}\}$, respectively. Each query token vector $q_{i}$ interacts with every document token vector in $\mathcal{D}$, and the contribution of the query token to the overall relevance score is defined by the maximum similarity across document tokens. The overall relevance score is then computed by summing these maximum contributions across all query tokens. Formally, the document’s overall score is computed as
$s(\mathcal{Q},\mathcal{D})=\sum_{i=1}^{m}\max_{j\in[1,n]}\text{sim}(q_{i},d_{j}),$ (1)
where $\text{sim}(q_{i},d_{j})$ denotes the cosine similarity between query token $q_{i}$ and document token $d_{j}$. Despite involving complex token-wise similarity computations, practical efficiency is supported by an efficient indexing following [santhanam2022colbertv2effectiveefficientretrieval].
4.2 Query Encoder
The query encoder comprises image and text encoders and a transformer block with a cross-attention layer, each paired with its own projection layers. Given a user query composed of an image and text $(I^{(q)},T^{(q)})$, the goal is to produce the query feature set $\mathcal{Q}$ for subsequent scoring. The textual and visual inputs are first processed separately by a ColBERTv2 text encoder [santhanam2022plaidefficientenginelate] and a CLIP image encoder [radford2021learning], yielding a set of $N_{t}$ token-level text embeddings $\mathcal{T}^{(q)}=\{t^{(q)}_{1},t^{(q)}_{2},\dots,t^{(q)}_{N_{t}}\}$ and a global image feature $g^{(q)}$, respectively. In addition, a set of patch-level features $\mathcal{P}^{(q)}=\{p^{(q)}_{1},p^{(q)}_{2},\dots,p^{(q)}_{N_{p}}\}$ is extracted from the penultimate layer of the CLIP encoder, where $N_{p}=H\times W$ denotes the number of spatial patches. These patch features are then contextualized using the textual features $\mathcal{T}^{(q)}$ through the transformer block with cross-attention [vaswani2023attentionneed]. Specifically, $\mathcal{P}^{(q)}$ serves as the queries, while $\mathcal{T}^{(q)}$ provides the keys and values, producing a set of $N_{p}$ multimodal features $\mathcal{M}^{(q)}=\{m^{(q)}_{1},m^{(q)}_{2},\dots,m^{(q)}_{N_{p}}\}$. Finally, the query feature set $\mathcal{Q}$ is formed by projecting the global image feature $g^{(q)}$, all textual features $t^{(q)}_{i}$, and all multimodal features $m^{(q)}_{j}$ into $\mathbb{R}^{128}$ using their dedicated projection layers. Denoting the projected features by $\widehat{g}^{(q)}$, $\widehat{\mathcal{T}}^{(q)}=\{\widehat{t}^{(q)}_{i}\}$, and $\widehat{\mathcal{M}}^{(q)}=\{\widehat{m}^{(q)}_{j}\}$, respectively, the final query feature set is $\mathcal{Q}=\{\widehat{g}^{(q)}\}\cup\widehat{\mathcal{T}}^{(q)}\cup\widehat{\mathcal{M}}^{(q)}$ where all elements lie in $\mathbb{R}^{128}$.
4.3 Document Encoder
Similarly to the query encoder, our objective is to obtain the document feature set $\mathcal{D}$. We assume that each document contains a text body and an image depicting the main entity, where the image is augmented following [deng-etal-2025-muka]. Given a collection of documents in KB, we further augment each document by collecting images of related entities mentioned within the document. We then extract sets of textual, visual, and multimodal features—analogous to the query encoding process—from both the document text body and the collected images, as detailed below.
Augmenting Documents with Images of Related Entities Given the textual content $T^{(d)}$ of a document, we first identify all named entities mentioned within $T^{(d)}$. For each extracted entity, we then collect the corresponding document from the KB, whose main entity matches that extracted entity. Since each document in the KB includes an image of its main entity, we extract these images and use them as related-entity images. By collecting these images, the augmented document becomes $(I^{(d)}_{0},I^{(d)}_{1},\dots,I^{(d)}_{R},T^{(d)})$, where $I^{(d)}_{0}$ is the image of the main entity and $I^{(d)}_{r}$ for $r\in\{1,\dots,R\}$ are the images of related entities. Note that while we rely on an LLM to identify related entities in order to maintain compatibility with the KB structure released in prior work [deng-etal-2025-muka], one could alternatively use the hyperlinks available within Wikipedia in a practical deployment.
Textual and Visual Feature Extraction Given a multi-image–augmented document $(I^{(d)}_{0},I^{(d)}_{1},\dots,I^{(d)}_{R},T^{(d)})$, we extract textual and visual features from $T^{(d)}$ and the main entity image $I^{(d)}_{0}$ using the same ColBERTv2 text encoder and CLIP image encoder employed in the query encoding stage. This yields a set of token-level textual features $\mathcal{T}^{(d)}=\{t^{(d)}_{1},t^{(d)}_{2},\dots,t^{(d)}_{N_{t}}\}$ and a global image feature $g^{(d)}_{0}$. In MuKA, the document embedding is formed by projecting and merging these features. However, since our augmented document includes multiple images, we additionally encode the related-entity images $I^{(d)}_{r}$ into their corresponding global features $g^{(d)}_{r}$ using the same CLIP encoder. This straightforwardly extends MuKA’s document embedding mechanism for multi-image setup without introducing any additional learnable parameters.
Multimodal Feature Extraction While the above multi-image extension is simple yet effective, the resulting document embeddings remain limited to unimodal features, treating text and images independently, missing information that can only be captured by jointly modeling both modalities. To incorporate such multimodal interactions, we additionally adopt a multimodal mapping layer—a transformer block with a cross-attention layer. Following the procedure used in query encoding, we extract additional patch-level visual features $\mathcal{P}^{(d)}_{r}=\{p^{(d)}_{r,1},p^{(d)}_{r,2},\dots,p^{(d)}_{r,N_{p}}\}$ from each image $I^{(d)}_{r}$, for both the main entity and related entities, i.e., $r\in\{0,\dots,R\}$. We then compute multimodal features by contextualizing these patch embeddings with the textual features. Specifically, $\mathcal{P}^{(d)}_{r}$ serves as the query inputs, and $\mathcal{T}^{(d)}$ provides the key–value inputs to the cross-attention layer, producing a set of multimodal features $\mathcal{M}^{(d)}_{r}=\{m^{(d)}_{r,1},m^{(d)}_{r,2},\dots,m^{(d)}_{r,N_{p}}\}$. Finally, the document feature set $\mathcal{D}$ is constructed by projecting all textual, visual, and multimodal features using their dedicated projection layers:
$\mathcal{D}=\widehat{\mathcal{T}}^{(d)}\cup\bigcup_{r=0}^{R}\left(\{\widehat{g}^{(d)}_{r}\}\cup\widehat{\mathcal{M}}^{(d)}_{r}\right)$. (2)
Here, ${\widehat{g}^{(d)}_{r}}\cup\widehat{\mathcal{M}}^{(d)}_{r}$ denotes the set containing the projected global visual feature and the projected multimodal features associated with the image $I^{(d)}_{r}$, and $\widehat{\mathcal{T}}^{(d)}$ is the set of projected textual features.
Table 1: Results on Multimodal KB-VQA Benchmarks. Retrieval and QA performance are reported. All models are jointly trained on EVQA and RETINA. The GT Document row represents an upper bound using gold knowledge. MIMIR consistently outperforms prior methods in both settings with and without visual shortcuts, showing the benefit of leveraging multi-image context over single-image or text-only approaches.
InfoSeek EVQA RETINA (Seen) RETINA (Unseen)
Method Document Modality Recall@5 Acc Recall@5 BEM Recall@5 BEM Recall@5 BEM
PreFLMR [lin2024preflmr] text 40.7 33.5 63.9 57.0 23.6 19.9 11.1 13.6
UniIR [wei2024uniir] text + single-image 34.5 26.4 22.0 25.7 13.8 14.6 4.7 11.2
ReT [caffagni2025recurrence] text + single-image 50.1 40.4 46.3 43.7 15.1 15.2 6.8 11.4
MuKA [deng-etal-2025-muka] text + single-image 51.7 41.0 64.9 60.7 20.8 18.7 8.6 12.2
MIMIR (Ours) text + multi-image 53.3 43.2 65.8 62.2 43.2 32.7 36.9 31.6
GT Document 76.5 88.2 68.5 70.2
Entity Token Embedding
The multimodal mapping layer described above mixes each image with the full textual context. However, in this formulation, the entire set of text-token embeddings $\mathcal{T}^{(d)}$ is used to contextualize the patch embeddings of every image, regardless of which tokens are actually semantically relevant to that image. As a result, the full burden of cross-modal alignment is placed on the implicit attention mapping, while the explicit and known associations between each image and its corresponding named entity in $\mathcal{T}^{(d)}$ are completely disregarded. To address this limitation, we introduce an Entity Token Embedding (ETE), denoted $\theta_{\text{ETE}}$}, a learnable vector added to the text tokens of the named entity corresponding to each image. Specifically, for each image $I^{(d)}_{r}$, we construct an image-specific textual feature set $\mathcal{T}^{(d)}_{r}$ by adding the ETE vector to the text tokens associated with its corresponding named entity. That is, we update $t^{(d)}_{s}\leftarrow t^{(d)}_{s}+\theta_{\text{ETE}}$} for all $s\in\mathcal{S}_{r}$, where $\mathcal{S}_{r}$ denotes the set of token indices representing the named entity linked to image $I^{(d)}_{r}$. Finally, the multimodal features $\mathcal{M}^{(d)}_{r}$ are extracted for each image by feeding the entity-specific textual embeddings $\mathcal{T}^{(d)}_{r}$ and the patch-level visual features $\mathcal{P}^{(d)}_{r}$ into the multimodal mapping layer. This design enables the network to more clearly distinguish the tokens corresponding to the target entity from the remaining text tokens during multimodal feature extraction, resulting in a more informed and semantically aligned fusion for each image. The subsequent processing follows the same procedure as above.
5 Experiments
5.1 Experimental Setup
Datasets We evaluate models on our RETINA dataset as well as two existing benchmarks, EVQA [mensink2023encyclopedic] and InfoSeek [chen2023can]. Since RETINA focuses on examples without visual shortcuts, whereas EVQA and InfoSeek include examples with visual shortcuts, the combination of these datasets enables a comprehensive assessment of realistic MKB-VQA scenarios. In terms of scale, RETINA contains 120K training samples and a 2K human-curated test set; EVQA provides 167K training samples and 3.7K test samples; and InfoSeek offers 676K training samples and 4.7K test samples. The knowledge bases of all three benchmarks consist of textual Wikipedia documents (details are available in Supp. Mat.) Following [deng-etal-2025-muka], we augment these documents with an image of their corresponding main entity. Building on this, we further enrich each document by attaching multiple images of related entities—identified as described in Sec. 4.3—thereby enabling multimodal retrieval with document embeddings augmented by multiple images.
Metrics We evaluate performance on both multimodal document retrieval and final answer generation. For retrieval, we adopt Recall@$K$, a standard metric for assessing retrieval systems, which measures the proportion of queries for which the annotated relevant document appears within the top-$K$ retrieved results. For answer generation, we use BERT Matching (BEM) [bulian-etal-2022-tomayto] on RETINA and EVQA, which scores a prediction by computing token-level matches weighted by contextual similarity from a BERT model, providing a soft semantic measure beyond strict lexical matching. For InfoSeek, we report the Accuracy metric specifically designed for this benchmark, which measures the proportion of correct answer predictions based on predefined matching functions tailored to each answer type [chen2023can]. All retrievers are evaluated by assessing the answers of the same VILA-13B [liu2023visualinstructiontuning] using their top-5 results following [deng-etal-2025-muka].
Models We compare our proposed MIMIR against four established multimodal baselines: UniIR [wei2024uniir], ReT [caffagni2025recurrence], PreFLMR [lin2024preflmr], and MuKA [deng-etal-2025-muka]. Among these, MIMIR, PreFLMR, and MuKA share largely the same core encoder architecture: they employ ColBERT-base-v2 for text encoding and CLIP ViT-G for visual encoding, along with their respective projection layers. The main difference lies in the design of the document encoder. PreFLMR operates on text-only documents, while MuKA enriches each document by attaching a single image corresponding to the main entity. In contrast, MIMIR further extends the document representation by incorporating multiple images of related entities within each document, as described in Sec. 4.3.
5.2 Results
Table 2: MIMIR Component Ablation. From top to bottom: MuKA [deng-etal-2025-muka]; extended MuKA with multi-image input; with multimodal features; and full MIMIR with entity token embeddings.
RETINA
# MI MMF ETE InfoSeek EVQA Seen Unseen
1 51.7 64.9 20.8 8.6
2 51.5 65.1 38.1 33.3
3 51.6 65.3 41.1 34.5
4 53.3 65.8 43.2 36.9
Comparisons to SOTA Multimodal Retrievers In Tab. 1, we compare MIMIR with state-of-the-art methods on InfoSeek, EVQA, and our RETINA dataset, where all models are jointly trained on the combined EVQA and RETINA training sets. Multimodal models equipped with a single image generally outperform text-only models on InfoSeek and EVQA; however, the trend reverses on RETINA, where text-only models achieve higher retrieval performance. This contrast highlights the adverse effects of visual shortcuts—multimodal models with a single image tend to overfit to shortcut cues and fail to effectively integrate visual and textual information when such cues are absent. Nevertheless, even the strongest baselines exhibit very limited performance on RETINA. Note that all models are trained on datasets containing both shortcut and non-shortcut samples. These results underscore the value of RETINA and validate the challenges it poses. In contrast, our proposed MIMIR achieves nearly twice the recall of the best-performing baselines and substantially improves answer quality, demonstrating the effectiveness of our document embeddings augmented with multiple images of related entities. Additionally, we report the answer generation performance using GT documents to provide an upper bound, and the results indicate that there is still considerable room for improvement.
Ablation Studies
Tab. 2 examines the use of multiple images (MI), multimodal feature (MMF), and entity token embeddings (ETE) discussed in Sec. 4.3. Without any of these components (Row 1), corresponding to the original MuKA model, performance on RETINA remains poor. Simply introducing multiple images (Row 2) leads to clear improvements on RETINA, while maintaining comparable performance on InfoSeek and EVQA. Since MuKA does not employ multimodal patch embeddings for document representations, this variant simply concatenates the projected global image features of the additional images. When multimodal patch embeddings are further incorporated for these augmented images (Row 3), performance improves even more, again without degradation on the existing benchmarks. Finally, adding ETE within the multimodal patch-embedding extraction process yields significant gains, as it allows cross-attention to focus more effectively on text tokens aligned with the image content.
(a) Seen
(b) Unseen
Figure 5: Analysis of Visual Shortcuts on RETINA. MIMIR (green) and MuKA (red) curves show recall for distractor (dotted) and GT (solid) documents. Distractors are non-GT documents whose main entity matches the query image’s entity.
Analysis of Visual Shortcuts We further analyze the results of MIMIR and MuKA, both trained and tested on RETINA. Specifically, we examine how these models are affected by non-target documents whose main entity corresponds to the query image, which we refer to as ”distractors”. In Fig. 5, we report recall@$k$ against the GT document (solid curves) and against the distractors (dotted curves). On the seen subset, at recall@1, MuKA retrieves distractors more frequently than GT documents. Although GT documents are retrieved more often than distractors as $k$ increases, the gap between the two recall curves remains narrow, suggesting that the model is easily confused by visually similar document images. This effect is pronounced in the unseen subset, where distractors exceed GT documents. In the 14.5% of cases where a distractor was retrieved, 78% of those cases contained only distractors within the top-5 results. This behavior stems from MuKA’s single-image augmentation strategy, which induces an incorrect inductive bias favoring visual shortcuts. In contrast, our MIMIR, through multi-image augmentation, mitigates this bias by incorporating diverse entity representations for each document. As a result, MIMIR consistently achieves higher recall for GT documents across both seen and unseen subsets, with large margins over distractors, demonstrating the effectiveness of our proposed multi-image–augmented document embedding.
Figure 6: Qualitative Comparison of Retrieval on the RETINA Benchmark. Results for (a) MuKA [deng-etal-2025-muka] and (b) MIMIR. GT documents are in green; distractors containing the query image entity are in red; query entity in GT documents are highlighted in green. While MuKA is biased toward visual shortcuts and often retrieves documents containing images similar to the query, MIMIR leverages multiple related image augmented document embeddings to correctly retrieve the GT document.
Figure 7: Further Analysis on Existing Benchmarks. (a) shows the query with an image of a Wood duck; (b) shows the GT document of Wood duck; (c) and (d) show the top-5 retrieved documents from MuKA and MIMIR, respectively. Related entities (green) and answers (red) are highlighted in the text.
Qualitative Analysis We analyze qualitative examples comparing MuKA [deng-etal-2025-muka] and MIMIR, examining how augmenting documents with multiple images of related entities affects performance on RETINA. MuKA tends to retrieve documents with images that are visually similar to the query, exhibiting a bias towards visual shortcut (e.g., trees or snowy landscapes in Fig. 6a). As a result, MuKA often fails to retrieve the GT document in RETINA, since the examples in our dataset are explicitly designed to eliminate such visual shortcuts. In contrast, MIMIR augments each document with multiple images of related entities, allowing the inclusion of an image that corresponds to the same entity depicted in the query, thereby enabling successful retrieval. As shown in Fig. 6b, related entities such as ‘Fraxinus americana’ and ‘Nunavut’ (marked in green) correspond to the entity of the query image, illustrating how this multi-image augmentation facilitates successful retrieval.
Further Analysis on Existing Benchmarks Interestingly, we observe consistent improvements over MuKA [deng-etal-2025-muka] on existing benchmarks that contain visual shortcuts, even though MuKA already augments each document with an image of its main entity. Our qualitative analysis reveals that even in benchmarks with visual shortcuts, the query image and the main-entity image in the GT document may differ visually (e.g., male vs. female wood ducks in Fig. 7a and 7b), which can lead to retrieval failures (e.g., MuKA results in Fig. 7c). In contrast, the proposed method further augments each document with images of related entities within the same GT document (e.g., entities highlighted in green in Fig. 7b), some of which are visually similar entities to the main entity (e.g., a mandarin duck in Fig. 7). This enrichment leads to a increased likelihood of retrieving the correct document. For example, Fig. 7d shows that the GT document for the wood duck is successfully retrieved because the document was augmented with the image of a mandarin duck—previously one of the most challenging negative samples for the baseline. This success is likely facilitated by structure of the Wikipedia, as related entities within it often depict visually similar entities.
6 Conclusion
We propose a novel MKB-VQA benchmark, RETINA, which reflects a challenging subset of real-world scenarios, without visual shortcuts. We show that previous MKB-VQA benchmarks allow models to exploit visual shortcuts, limiting their practical utility. To address this, we introduce MIMIR, a retriever that augments multiple images of related entities to the document embedding, effectively solving RETINA. Looking ahead, real-world scenarios involve diverse modalities of related context beyond images, and extending our approach to incorporate these modalities presents a promising direction.
[…]
Supplementary Material
Appendix A RETINA Bench Construction
This section describes the process of constructing RETINA bench, including how textual queries and images were collected, processed, and curated to ensure quality. Throughout all stages of this pipeline, Qwen2.5-32B [bai2023qwen] was used.
A.1 Query and Document Image Pool Construction
In this subsection, we explain how the pools of question and document images were created, including data sources, replacement strategies for unavailable images, and verification procedures to avoid duplication. Images for questions and documents were sourced from separate pools. Question images were taken from the M2KR dataset, including OVEN [hu2023opendomainvisualentityrecognition], Google Landmark v2 [weyand2020googlelandmarksdatasetv2], and iNaturalist [vanhorn2018inaturalistspeciesclassificationdetection]. Document images were downloaded from web page links provided by MuKA [deng-etal-2025-muka]. For images that were no longer available, replacements were retrieved using the Wikipedia API and Google Lens, and verified with the imagehash toolkit to ensure they were not identical to the query images, we used a perceptual hash size of 8, treating a distance of 10 or less as identical following [deng-etal-2025-muka].
A.2 Paraphrasing
Here, we describe the paraphrasing process applied to the questions to prevent them from being directly retrievable using the text alone. We evaluate retrieval using BM25 [robertson1995okapi] with a top-5 setting. Before paraphrasing, approximately 37% of the noisy samples are retrievable. After paraphrasing, recall drops from 37% to 25%, and the remaining retrievable samples are removed from the dataset.
A.3 Human Curation
Figure 8: Human Curation User Interface An image, document title, a question, and an answer are presented for curation, along with guidelines for identifying common issues.
Figure 9: Qualitative Examples of RETINA.
Figure 10: Filtered Examples with Multiple Answers.
Figure 11: Additional Examples of MIMIR retrieved documents compared to MuKA.
This subsection details the human curation procedure, including the criteria for filtering questions, the curation workflow, and statistics on how many questions were retained versus excluded. Three fluent english-speaking computer science students participated. During curation, questions that could have multiple valid answers, questions answerable without viewing the associated image, questions containing the answer within the question itself Fig. 8, grammatically awkward questions, and any other potentially problematic questions were filtered out. A total of 6k randomly sampled questions were examined, resulting in 1k curated test samples each for seen and unseen set. The curators followed a shared written guideline throughout the process, which took roughly 24 hours in total. Following this process, 29.16% of the questions passed curation and were considered high-quality, while the remaining questions were excluded. Problematic questions were further categorized according to common issues: 21.27% were solvable without the image (solvable_without_image), 75.35% had multiple correct answers (multiple_answers), 0.35% contained grammatical errors (grammar_issue), 2.26% included lexical cues (lexical_cues), 0.15% had images that explicitly revealed the answer (explicit_image_cue).
Appendix B RETINA Qualitative Examples
In this section, we present qualitative examples from RETINA. As shown in Figure 9, the test set consists of samples selected through human curation to ensure natural, high-quality instances. In contrast, Figure 10 shows samples that were removed during curation because they contain multiple plausible answers (75.35% of the filtered samples). These ambiguous cases remain in the raw training set, where they are still useful for learning, but are excluded from the test set because they make evaluation unreliable.
Table 3: Standard Deviation Analysis. Average retrieval performance and standard deviations of MIMIR are reported.
Method Document Modality InfoSeek EVQA RETINA (Seen) RETINA (Unseen)
PreFLMR [lin2024preflmr] text 40.7 63.9 23.6 11.1
UniIR [wei2024uniir] text + single-image 34.5 22.0 13.8 4.7
ReT [caffagni2025recurrence] text + single-image 50.1 46.3 15.1 6.8
MuKA [deng-etal-2025-muka] text + single-image 51.7 64.9 20.8 8.6
MIMIR (Ours) text + multi-image 53.1 ($\pm 0.5$) 65.5 ($\pm 0.4$) 43.0 ($\pm 0.7$) 36.5 ($\pm 0.8$)
Appendix C MIMIR Qualitative Examples
In this section, we illustrate qualitative examples of MuKA[deng-etal-2025-muka] (Fig. 11a) and MIMIR (Fig. 11b). Because RETINA is designed to avoid visual shortcuts, MuKA’s tendency to retrieve documents with images that are visually similar to the query leads it to select incorrect documents. In contrast, MIMIR augments the document with multiple images of related entities, allowing a visual match with the query. This visual evidence reinforces the textual relevance of the GT document, enabling successful retrieval.
Appendix D KB Construction for Inference
In this section, we describe how the KB is constructed for retrieval, including the passages used and their configurations. EVQA and RETINA share the same KB, built from the EVQA subset of Wikipedia passages provided by the M2KR benchmark, while Infoseek uses a separate KB constructed from the Infoseek subset of M2KR Wikipedia passages [lin2024preflmr]. Since RETINA shares the same KB as EVQA, only samples whose GT documents do not appear in the EVQA training set are included as unseen test samples for RETINA. For KB indexing with PLAID [santhanam2022plaidefficientenginelate], we set the parameters as follows: the number of bits to 8, the number of k-means iterations to 20, used cosine similarity and the batch size was set to 64 following [santhanam2022colbertv2effectiveefficientretrieval, deng-etal-2025-muka, lin2024preflmr].
Appendix E Implementation Details
In this section, we provide implementation details, including the dataset construction process and model configurations. Following MuKA [deng-etal-2025-muka], the model is initialized with the weights of PreFLMR [lin2024preflmr] using the CLIP ViT-G variant, which has demonstrated the best performance. Training was conducted for one epoch using the Adam optimizer. The projection layers are 2 layer MLP following [deng-etal-2025-muka, lin2024preflmr] for projecting the multimodal features as 32 tokens embeddings and lowering the dimension to 128 for efficiency [lin2023finegrainedlateinteractionmultimodalretrieval, lin2024preflmr, deng-etal-2025-muka]. For the number of augmented related images, we do not impose an upper bound on the number of augmented related images, which is an average 4.3 per sample. For training, we utilize 120k training sample randomly sampled from EVQA and RETINA, respectively. For the preliminary experiment, we discarded the text token features from the query embeddings while keeping all other architectural components and settings identical [deng-etal-2025-muka]. For answer generation, the VILA 13B model [lin2024vilapretrainingvisuallanguage] employed in [deng-etal-2025-muka] was used. The top-5 results from the zero-shot inference of the PreFLMR [lin2024preflmr] were utilized to perform LoRA training on VILA 13B. LoRA rank and alpha were set to 128 and 256, respectively, with a batch of 512. The Adam optimizer was applied, training only the multimodal projectors and LoRA modules with learning rates of 2e-5 and 2e-4 [deng-etal-2025-muka].
VILA Training Prompt
<image>
Question: question
Retrieved passages:
1: <image>document_text_1
2: <image>document_text_2
3: <image>document_text_3
4: <image>document_text_4
5: <image>document_text_5
Given the query image and question,
identify matched passages and provide a short answer.
Figure 12: VILA Training Prompt.
Retriever Instruction Prompts
 <Image> Using the provided image, obtain
documents that address the subsequent question:
<Image> Retrieve documents that provide an
answer to the question alongside the image:
<Image> Extract documents linked to the
question provided in conjunction with the image:
<Image> Utilizing the given image, obtain
documents that respond to the following question:
<Image> Using the given image, access documents that
provide insights into the following question:
<Image> Obtain documents that correspond to
the inquiry alongside the provided image:
<Image> With the provided image, gather documents
that offer a solution to the question:
<Image> Utilizing the given image, obtain
documents that respond to the following question:
Figure 13: Retriever Instruction Prompts.
Appendix F Standard Deviation Analysis
This section analyzes how consistent MIMIR’s performance is across multiple training runs. We trained the model five times with different random seeds. The resulting standard deviation across datasets was 0.6 on average.
Appendix G Limitations
MIMIR incorporates additional image features into document embeddings, increasing the similarity computations during retrieval. This leads to higher inference latency compared to using embeddings with a single image.
Appendix H Prompts
In this section, we provide the prompt templates used throughout our experiments.
Entity extraction prompt
You are a Named Entity Recognition and Relation Extraction system specialized for identifying entities that
correspond to Wikipedia article title.
Your task is to extract all distinct named entities mentioned or referenced in the passage that align with
Wikipedia article titles, even if they are not exact substrings.
Guidelines:
1. For each entity, return an object with the fields: "entity", "entity_type", and "relation".
2. The "entity" must be a canonical Wikipedia title or a close variant that accurately represents the mention
in the passage.
3. "entity_type" describes the category of the entity (e.g., Person, Place, Organization, Event, Book, Law, etc.).
You may create new types as needed.
4. The "relation" must follow these strict rules:
   - It must be a single, full natural language sentence.
   - The sentence must explicitly include the main entity as the subject.
   - It must describe concrete and fact-based relationships to the given entity, including specific information
   such as dates, functions, locations, roles, historical context, or other verifiable details.
   - If there are multiple distinct factual relations for the entity, combine them into one coherent
   and grammatically correct sentence.
   - The sentence must be informative enough to support downstream question generation tasks.
   - Avoid abstract, vague, or generic relations (e.g., "is related to", "is part of", or incomplete phrases).
5. If the passage does not provide any clear relation for the entity, set "relation" to null.
6. Merge variations or aliases of the same entity to avoid duplication.
7. Be comprehensive and tolerant of indirect mentions or paraphrased references.
Wikipedia Title: "{entity}"
Passage:
Figure 14: Entity extraction prompt.
Question Generation Prompt
You are a VQA question-answer pair generator.
Given:
- Main entity (answer to the question)
- Related entity (depicted in the image)
- Relation sentence describing how main and related entities are connected
- Attribute sentence describing the main entity
Generate a single, grammatically correct question-answer pair satisfying:
- The question refers indirectly to the related entity shown in the image by including exactly one singular
demonstrative phrase matching the related entity’s type, e.g. "this country", "this building", "this city", etc.
- Do NOT explicitly name the related entity in the question.
- The demonstrative phrase must appear exactly once and without any article (no "a", "an", or "the").
- The question must use only the relation and/or attribute sentences to describe the main entity indirectly.
- The answer must be exactly the main entity’s name.
- Do NOT mention or describe the image explicitly (no "shown here", "in this image", etc.).
- Output the question and answer in the format:
Question: [your question]
Answer: [main entity]
Figure 15: Question Generation Prompt.
Paraphrasing Prompt
You are a helpful assistant that rewrites trivia questions only.
Keep the original meaning and difficulty but change the wording.
Do NOT include the answer in your response — output ONLY the rewritten question.
Rewrite the following location-based trivia question using different wording,
but keep the original meaning and difficulty level.
Output format:
Question: [your rewritten question]
Figure 16: Paraphrasing Prompt.